# Transitivity of the pure Hurwitz classes of quadratic post-critically finite polynomials

### v1

@@ Abstract.
We prove that for two post-critically finite quadratic polynomials $f,g$, there is a mapping class $\phi$ of the sphere with finitely many marked points such that $f\phi$ and $g$ are pure Hurwitz equivalent.
@@ 1. Introduction
The study of the dynamics of post-critically finite rational maps dates back to the work of Fatou and Julia in the 1910s and has gained traction since the work of Thurston in the 1980s. Moreover, the pure mapping class group acts on the set of post-critically finite branched covers with a given post-critical set and preservers the branching data of the maps. This action has been of considerable interest as highlighted in work of Bartholdi–Dudko, Bartholdi–Nekrashevych, Kelsey–Lodge, Koch–Pilgrim–Selinger, Pilgrim, D. Thurston (see [1, 2, 3, 5, 6, 7, 8, 9, 10], for example), and connects holomorphic dynamics, geometric group theory and surface topology. On the other hand, impure mapping class groups also act on the set of post-critically finite branched covers, but may change the combinatorics of the branched cover. This action is a priori richer, but requires an explicit understanding of how impure mapping classes change the combinatorics of branching data. In this paper, we make a first step in this direction, showing that for any two pure Hurwitz classes – a combinatorial description of the dynamics of the branched cover – there is an impure mapping class between representatives of the classes in the quadratic case.
@@
Combinatorial and portrait equivalence. In the study of holomorphic dynamics, post-critically finite marked branched covers are often considered up to combinatorial (or Thurston) equivalence. That is: let $f$ and $g$ be post-critically finite marked branched covers with marked sets $M_{f}$ and $M_{g}$, respectively. We say that $f$ and $g$ are combinatorially equivalent if there exist orientation preserving homeomorphisms $h_{1},h_{2}$ such that:
@@
    $h_{1}$ and $h_{2}$ are homotopic relative to $M_{f}$.
The branched covers $f,g$ are pure Hurwitz equivalent if the first two conditions above are satified and $h_{0}(p)=h_{1}(p)$ for all $p\in P_{f}$. Kelsey–Lodge prove that two post-critically finite quadratic rational maps are pure Hurwitz equivalent if and only if their dynamical portraits are isomorphic as directed graphs [7, Lemma 3.6]. Our main theorem is that given any two pure Hurwitz classes, there is a mapping class that takes a representative of one to a representative of the other.
@@
Let $f,g:S^{2}\rightarrow S^{2}$ be marked quadratic branched covers with dynamical portraits $\Gamma_{f}$ and $\Gamma_{g}$. There exists a finite set of points $M\subset S^{2}$ that contains the post-critical set for $f$ and $g$ and $\phi\in\operatorname{Mod}(S^{2},M)$ so that $f\circ\phi$ is pure Hurwitz equivalent to $\Gamma_{g}$.
In particular, $f$ and $g$ may have different numbers of post-critical points. Using an impure mapping class $\phi$ allows us to increase or decrease the number of post-critical points of the branched cover.
@@
Strategy. We prove Theorem 1.1 by developing an algorithm that finds a sequence of marked branched covers $\{(f,M_{f})=(f_{0},M_{0}),(f_{1},M_{1}),\cdots,(f_{n},M_{n})=(g,M_{g})\}$ such that each $f_{i}$ differs from $f_{i+1}$ by what we call a (portrait) function. Most functions are half twists either between points in $M_{i}$ or between one point in $M_{i}$ and a point that is either a fixed point of $f_{i}$ or is in $f^{-1}(M_{i})$. If the function by which we compose is a half twist between a point in $M_{i}$ and another point $\star$, we take $M_{i+1}=M_{i}\cup\{\star\}$. Some functions are a composition of two such half-twists.
The overall strategy is as follows: start with the marked branched cover $(f,M_{f})=(f_{0},M_{0})$. We create a sequence of marked branched covers $\{(f_{i},M_{i})\}$ where the portrait for $f_{i+1}$ differs from the portrait from $f_{i}$ by one of the functions 1-13 defined in Section 3. This process terminates when we reach a branched cover $f_{n}$ that has the same portrait as $g$. For each $i\geq 1$, the algorithm compares the portrait of $f_{i}$ to the portrait of $g$, and determines a mapping class by which we pre-compose $f_{i}$ to obtain the next map $f_{i+1}$.
Marked set. Theorem 1.1 states that marked branched covers $(f,M_{f})$ and $(g,M_{g})$ that for some marked set $M$, there is a mapping class $\phi\in\operatorname{Mod}(S^{2},M)$ for which the portrait of $f\phi$ is isomorphic to the portrait of $g$. As indicated above, the algorithm determines this set by first beginning with the post-critical set of $M_{0}=M_{f}$ and then adding additional points at some stages of the algorithm. At the $i$th stage, we add at most one point to the set $M_{i-1}$. The point that is (possibly) added is either a fixed point of $f_{i-1}$ or a preimage of a point that is either critical or post-critical under $f_{i}$. Because the algorithm terminates in a finite number of steps, the set $M$ consists of the union of $M_{f}$ with a finite set of points.
@@ 1.1. Future directions
Much of the study of actions of mapping class groups on sets of post-critically finite branched covers has restricted to actions of pure mapping class groups. The primary goal of this paper is to promote the study of the action of impure mapping class groups on post-critically finite branched covers.
Moreover, the functions we present in Section 3 to change dynamical portraits provide a tool to use arguments that induct on the number of points of a post-critical set in families of post-critically finite branched covers.
Braids versus symmetric groups. As indicated above, we construct a sequence of marked branched covers $\{(f,M)=(f_{0},M_{0}),(f_{1},M_{1}),\cdots,(f_{n},M_{n})=(g,M_{g})\}$, and keep track of a sequence of mapping classes $\{\phi_{1},\cdots,\phi_{n}\}$ for which $f_{i+1}=f_{i}\phi_{i+1}$. The functions are agnostic to the choice of mapping class up to its action on the marked set $M_{i}$. We show that we can “move between” different pure Hurwitz equivalence classes by composing by an impure mapping class. With additional information about the actual mapping class (rather than just its quotient to a finite symmetric group), it is possible that one can make stronger statements about finer equivalence classes (such as combinatorial equivalence).
@@ 1.3. Acknowledgments
The authors would like to thank ICERM where the project originated. The authors would like to thank Russell Lodge and Kevin Walsh for conversations about the paper and Dan Margalit for comments on an earlier draft of the paper. The authors would like to thank Kevin Pilgrim and the University of Indiana for hosting the authors at the Early-Career Complex Dynamics workshop in summer 2024 and Giulio Tiozzo for inviting the second author to University of Toronto. The first author was partially supported by an NSERC PDF, and was partially supported by NSERC Discovery Grant RGPIN-2024-05587. The second author was supported by the National Science Foundation under Grant No. DMS-2002951.
@@ 2.1. Abstract portraits
Abstract portraits for marked branched covers. Let $f$ be a marked post-critically finite branched cover. There is an abstract portrait associated to $f$ in which the vertices correspond to a subset of the marked set $M_{f}$. The vertices corresponding to critical points that are labeled $\ast$. Let $p_{1}$ and $p_{2}$ be points in $M_{f}$, and let $v_{i}$ be the vertex of the portrait of $f$ corresponding to $p_{i}$. There is an edge of the portrait from $v_{1}$ to $v_{2}$ if $f(p_{1})=p_{2}$. The edge will be labeled with the local degree of $f$ at $p_{1}$. The label will be greater than one if and only if $v_{1}$ is a critical point of $f$. We will suppress the labels when their value is one, or if they are not relevant to the discussion. We observe that because $f$ is a branched cover, the sum of the labels of the arrows that terminate at a vertex cannot exceed the degree of the branched cover. We can think of $f$ as acting on its dynamical portrait by mapping a vertex $v$ of the portrait of $f$ to the vertex at the terminus of the edge originating at $v$. In this way, we will think of the vertices of a portrait as the orbit of $f$.
@@ 2.2. Portrait features
Number of pre-periods. As above, any point contained in the cycle of a graph is periodic. If a point distinguished as a critical point is not periodic, there must be a directed path from it to (at least one) periodic point and we say that it is (strictly) pre-periodic. If a point distinguished as a critical point $C_{j}$ for $j\in\{1,2\}$ is pre-periodic, there is a (unique) shortest path from $C_{j}$ to a cycle of the graph $\{C_{j},P_{1},\cdots,P_{r}\}$ where $P_{r}$ is periodic and no other point in the path is periodic. We call the set of points $\{C_{j},P_{1},\cdots,P_{r-1}\}$ the pre-period of $C_{j}$.
Because every pre-period must begin with a point distinguished as a critical point, $\Gamma$ can have either zero, one, or two pre-periods (at most one for each point distinguished as critical points that it contains). If $\Gamma$ has two components, each component contains one point distinguished as a critical point. Each component has one pre-period for each pre-periodic critical point, so each component can have zero, one, or two pre-periods.
Pre-period patterns. If $\Gamma$ has one component and both points of $C$ are strictly pre-periodic, their pre-periods can either be disjoint or can intersect non-trivially.
If the pre-periods of $C_{1}$ and $C_{2}$ intersect non-trivially, there are two additional possibilities:
@@
If a component of $\Gamma$ contains both vertices in $C$ and one is periodic and the other is pre-periodic, then $\Gamma$ has one pre-period. Label the periodic point of $C$ as $C_{2}$ and the strictly pre-periodic point of $C$ as $C_{1}$. In particular, there will be a directed path from $C_{1}$ to $C_{2}$ but not from $C_{2}$ to $C_{1}$. Let $q$ be the length of the (shortest) directed path from $C_{1}$ to $C_{2}$ and let $r$ be the length of the pre-period of $C_{1}$ as above. We refer to $q-r$ as the position of $C_{2}$ in the cycle of $C_{1}$. We can also measure the length of the shortest path from $C_{2}$ to the first periodic point, which will be $k-(q-r)$.
@@
When the portrait has one component, there are five different pre-period structures. If both critical points are strictly pre-periodic, and their pre-periods intersect non-trivially, we need only measure the length of the cycle. In the remaining three cases, we observe that there are two points in the cycle of $\Gamma$ between which we measure distance:
@@ 3. Portrait functions
Throughout this section, let $f$ be a marked post-critically finite quadratic branched cover with portrait $\Gamma_{f}$ and marked set $M_{f}$. We will denote the critical points of $f$ by $C_{1}$ and $C_{2}$. As discussed above, one critical point may be in the $f$-forward orbit of the other. If this is the case, we assume (up to relabeling) that $C_{2}=f^{q}(C_{1})$ for some $q$. Let $P_{f}$ be the post-critical set of $f$ and let $N_{f}$ be the union of the fixed points of $f$ and $f^{-1}(P_{f}\cup\{C_{1},C_{2}\})$. Let $\phi\in\operatorname{Mod}(S^{2},N_{f})$. The map $f\phi$ will modify the portrait of $\Gamma_{f}$ and will have the property that $f(\phi(N_{f}))\subset N_{f}$. We call this modification a (portrait) function. The list below describes the portrait functions that we use by their desired effect on the portrait, though some of the functions require specific features of the starting portrait.
@@
    Increase a cycle length
  2. (2)
    Decrease a cycle length
  3. (3)
    Increase a pre-period length
  4. (4)
    Decrease a pre-period length
  5. (5)
    Move $C_{2}$ within the cycle of $C_{1}$ (if $C_{2}$ is in the $f$-orbit of $C_{1}$)
@@
    Make a strictly pre-periodic critical point periodic (if at least one critical point is pre-periodic)
@@
    Merge two disjoint components
@@
    Create a new pre-period from one component (if $C_{2}$ is not in the $f$-orbit of $C_{1}$)
@@
    Create a new pre-period from two components

### v2

@@ Abstract.
We prove that for two post-critically finite quadratic polynomials $g,h$, there is a mapping class $\phi$ of the sphere with finitely many marked points such that $g\phi$ and $h$ are pure Hurwitz equivalent.
@@ 1. Introduction
The study of the dynamics of post-critically finite rational maps dates back to the work of Fatou and Julia in the 1910s and has gained traction since the work of Thurston in the 1980s. Moreover, the pure mapping class group acts on the set of post-critically finite branched covers with a given post-critical set and preservers the branching data of the maps. This action has been of considerable interest by Bartholdi–Dudko, Bartholdi–Nekrashevych, Pilgrim, Kelsey–Lodge, D. Thurston, Koch–Pilgrim–Selinger (see [1, 2, 3, 6, 7, 8, 9, 10, 11], for example), and connects holomorphic dynamics, geometric group theory and surface topology. On the other hand, impure mapping class groups also act on the set of post-critically finite branched covers, but may change the combinatorics of the post-critical set of the branched cover. This action is a priori richer, but requires an explicit understanding of how impure mapping classes change the combinatorics of branching data. In this paper, we make a first step in this direction, showing that for any two pure Hurwitz classes – a combinatorial description of the dynamics of the branched cover – there is an impure mapping class between representatives of the classes in the quadratic case.
@@
Combinatorial and portrait equivalence. In the study of holomorphic dynamics, post-critically finite marked branched covers are often considered up to combinatorial (or Thurston) equivalence. That is: let $g$ and $h$ be post-critically finite marked branched covers with marked sets $M_{g}$ and $M_{h}$, respectively. We say that $g$ and $h$ are combinatorially equivalent if there exist orientation preserving homeomorphisms $\psi_{1},\psi_{2}$ such that:
@@
    $\psi_{1}$ and $\psi_{2}$ are homotopic relative to $M_{g}$.
The branched covers $g,h$ are pure Hurwitz equivalent if the first two conditions above are satified and $\psi_{0}(p)=\psi_{1}(p)$ for all $p\in P_{g}$. Kelsey–Lodge prove that two post-critically finite quadratic rational maps are pure Hurwitz equivalent if and only if their dynamical portraits are isomorphic as directed graphs [8, Lemma 3.6]. Our main theorem is that given any two pure Hurwitz classes, there is a mapping class that takes a representative of one to a representative of the other.
@@
Let $g,h:S^{2}\rightarrow S^{2}$ be marked quadratic branched covers with dynamical portraits $\Gamma_{g}$ and $\Gamma_{h}$. There exists a finite set of points $M\subset S^{2}$ that contains the post-critical set for $g$ and $h$ and $\phi\in\operatorname{Mod}(S^{2},M)$ so that $g\circ\phi$ is pure Hurwitz equivalent to $h$.
In particular, $g$ and $h$ may have different numbers of post-critical points. Using an impure mapping class $\phi$ allows us to increase or decrease the number of post-critical points of the branched cover.
@@
Strategy. We prove Theorem 1.1 by developing an algorithm that finds a sequence of marked branched covers $\{(g,M_{g})=(f_{0},M_{0}),(f_{1},M_{1}),\cdots,(f_{n},M_{n})=(h,M_{h})\}$ such that each $f_{i}$ differs from $f_{i+1}$ by what we call a (portrait) function. The functions are half twists either between points in $M_{i}$ or between one point in $M_{i}$ and a preimage of a critical point. If the function by which we compose is a half twist between a point in $M_{i}$ and another point $\diamond$, we take $M_{i+1}=M_{i}\cup\{\diamond\}$.
The overall strategy is as follows: start with the marked branched covers $(g,M_{g})=(f_{0},M_{0})$ and $(h,M_{h})$. We create sequences of marked branched covers $\{(f_{i},M_{i})\}$ where the portrait for $f_{i+1}$ differs from the portrait from $f_{i}$ by one of the functions 1-3 defined in Section 3. This process terminates when we reach a branched cover $f_{n}$ that has the same portrait as $z^{2}$. We repeat the process for $(h,M_{h})$. Then invert the sequence from $(h,M_{h})$ to $z^{2}$ to create a sequence from $(g,\Gamma_{g})$ to $(h,\Gamma_{h})$.
Marked set. Theorem 1.1 states that marked branched covers $(g,M_{g})$ and $(h,M_{h})$ that for some marked set $M$, there is a mapping class $\phi\in\operatorname{Mod}(S^{2},M)$ for which the portrait of $g\phi$ is isomorphic to the portrait of $h$. As indicated above, the algorithm determines this set by first beginning with the post-critical set of $M_{0}=M_{f}$ and then adding additional points at some stages of the algorithm. At the $i$th stage, we add at most one point to the set $M_{i-1}$. The point that is (possibly) added is a preimage of a point that is a critical point of $f_{i}$. We repeat the algorithm with $h$ to obtain another set of points consisting of $M_{h}$ and a union of points obtained in each step. Because the algorithm terminates in a finite number of steps, the set $M$ consists of the union of $M_{g}\cup M_{h}$ with a finite set of points.
@@ 1.1. Future directions
Much of the study of actions of mapping class groups on sets of post-critically finite branched covers has restricted to actions of pure mapping class groups. The primary goal of this paper is to promote the study of the action of impure mapping class groups on post-critically finite branched covers. Moreover, the functions we present in Section 3 to change dynamical portraits provide a tool to use arguments that induct on the number of points of a post-critical set in families of post-critically finite branched covers.
Braids versus symmetric groups. As indicated above, we construct a sequence of marked branched covers $\{(g,M_{g})=(f_{0},M_{0}),(f_{1},M_{1}),\cdots,(f_{n},M_{n})=(h,M_{h})\}$, and keep track of a sequence of mapping classes $\{\phi_{1},\cdots,\phi_{n}\}$ for which $f_{i+1}=f_{i}\phi_{i+1}$. The functions are agnostic to the choice of mapping class up to its action on the marked set $M_{i}$. We show that we can “move between” different pure Hurwitz equivalence classes by composing by an impure mapping class. With additional information about the actual mapping class (rather than just its quotient to a finite symmetric group), it is possible that one can make stronger statements about finer equivalence classes (such as combinatorial equivalence).
@@ 1.3. Acknowledgments
The authors would like to thank ICERM where the project originated. The authors would like to thank Russell Lodge, Dan Margalit, Kevin Pilgrim, and Kevin Walsh for conversations about the paper. The authors would like to thank Kevin Pilgrim and the University of Indiana for hosting the authors at the Early-Career Complex Dynamics workshop in summer 2024 and 2025 and Giulio Tiozzo for inviting the second author to University of Toronto. The first author was partially supported by an NSERC PDF, and was partially supported by NSERC Discovery Grant RGPIN-2024-05587. The second author was supported by the National Science Foundation under Grant No. DMS-2002951.
@@ 2.1. Abstract portraits
Abstract portraits for marked branched covers. Let $f$ be a marked post-critically finite branched cover. There is an abstract portrait associated to $f$ in which the vertices correspond to a subset of the marked set $M_{f}$. The vertices corresponding to critical points are labeled $\ast$. Let $p_{1}$ and $p_{2}$ be points in $M_{f}$, and let $v_{i}$ be the vertex of the portrait of $f$ corresponding to $p_{i}$. There is an edge of the portrait from $v_{1}$ to $v_{2}$ if $f(p_{1})=p_{2}$. The edge will be labeled with the local degree of $f$ at $p_{1}$. The label will be greater than one if and only if $v_{1}$ is a critical point of $f$. We will suppress the labels when their value is one, or if they are not relevant to the discussion. We observe that because $f$ is a branched cover, the sum of the labels of the arrows that terminate at a vertex cannot exceed the degree of the branched cover. We can think of $f$ as acting on its dynamical portrait by mapping a vertex $v$ of the portrait of $f$ to the vertex at the terminus of the edge originating at $v$. In this way, we will think of the vertices of a portrait as the orbit of $f$.
@@ 2.2. Portrait features
Number of pre-periods. As above, any point contained in the cycle of a graph is periodic. If a point distinguished as a critical point is not periodic, there must be a directed path from it to (at least one) periodic point and we say that it is (strictly) pre-periodic. If a point distinguished as a critical point $C_{j}$ for $j\in\{1,2\}$ is pre-periodic, there is a (unique) shortest path from $C_{j}$ to a cycle of the graph $\{C_{j},P_{1},\cdots,P_{r}\}$ where $P_{r}$ is periodic and no other point in the path is periodic. We call the set of points $\{C_{j},P_{1},\cdots,P_{r-1}\}$ the pre-period of $C_{j}$. Because every pre-period must begin with a point distinguished as a critical point, $\Gamma$ can have either zero, one, or two pre-periods (at most one for each point distinguished as critical points that it contains). If $\Gamma$ has two components, each component contains one point distinguished as a critical point. Each component has one pre-period for each pre-periodic critical point, so each component can have zero, one, or two pre-periods.
Pre-period patterns. If $\Gamma$ has one component and both points of $C$ are strictly pre-periodic, their pre-periods can either be disjoint or can intersect non-trivially. If the pre-periods of $C_{1}$ and $C_{2}$ intersect non-trivially, there are two additional possibilities:
@@
If a component of $\Gamma$ contains both vertices in $C$ and one is periodic and the other is pre-periodic, then $\Gamma$ has one pre-period. Label the periodic point of $C$ as $C_{2}$ and the strictly pre-periodic point of $C$ as $C_{1}$. In particular, there will be a directed path from $C_{1}$ to $C_{2}$ but not from $C_{2}$ to $C_{1}$. Let $q$ be the length of the (shortest) directed path from $C_{1}$ to $C_{2}$ and let $r$ be the length of the pre-period of $C_{1}$ as above. We refer to qrq-r as the position of $C_{2}$ in the cycle of $C_{1}$. We can also measure the length of the shortest path from $C_{2}$ to the first periodic point, which will be k(qr)k-(q-r).
@@
When the portrait has one component, there are five possible pre-period structures. If both critical points are strictly pre-periodic, and their pre-periods intersect non-trivially, we need only measure the length of the cycle. In the remaining three cases, we observe that there are two points in the cycle of $\Gamma$ between which we measure distance:
@@ 3. Portrait functions
Throughout this section, let $f$ be a marked post-critically finite quadratic branched cover with portrait $\Gamma_{f}$ and marked set $M_{f}$. We will denote the critical points of $f$ by $C_{1}$ and $C_{2}$. As discussed above, one critical point may be in the $f$-forward orbit of the other. If this is the case, we assume (up to relabeling) that $C_{2}=f^{q}(C_{1})$ for some $q$. Let $P_{f}$ be the post-critical set of $f$ and let $N_{f}$ be the union of $P_{f}$, $C_{1}\cup C_{2}$, and $f^{-1}(\{C_{1},C_{2}\})$. Let $\phi\in\operatorname{Mod}(S^{2},N_{f})$. The map $f\phi$ will modify the portrait of $\Gamma_{f}$ and will have the property that $f(\phi(N_{f}))\subset N_{f}$. We call this modification a (portrait) function. The list below describes the portrait functions that we use by their desired effect on the portrait, though some of the functions require specific features of the starting portrait.
@@
    Split the portrait from one to two components
@@
    Decrease a cycle length
For each function, we specify by which element $\phi$ of $\operatorname{Mod}(S^{2},N_{f})$ we pre-compose $f$ by to modify the portrait as desired. We then prove that the proposed composition modifies the portrait as desired. In the algorithm below, we apply the functions to target the portrait for $z^{2}$, which consists of two (disjoint) cycles of length one. To this end, we apply the functions in order; that is, we first apply Function 1 to a portrait, then, when applying Functions 2 and 3, we may assume the critical points are in distinct components. When applying Function 3, we may assume that the portrait has two cyclic components.
Function 1 (Split portrait from one to two components).
@@
    1. (a)
      If either $f(C_{1})=C_{2}$ or $f(C_{2})=C_{1}$, let $\phi$ be a half twist that swaps $C_{1}$ with $C_{2}$. Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ has two components, one where one critical point maps to itself, and the other critical point is in a cycle of period k1k-1.
    2. (b)
      Otherwise, let $q$ be such that $f^{q}(C_{1})=C_{2}$. Let $\phi$ be a half twist that swaps $f^{q-1}(C_{1})$ with $f^{k-1}(C_{1})$ (where $k$ is the length of the cycle). Output the marked branched cover $f\phi$, the marked set $M$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ has two cyclic components – $C_{1}$ is in a cycle of length $q$, $C_{2}$ is in a cycle of length kqk-q.
@@
    Suppose one critical point is in the pre-period of the other. Let $C_{2}$ be in the orbit of $C_{1}$, and let $q$ be such that $f^{q}(C_{1})=C_{2}$. Let $\diamond$ denote a pre-image of $C_{1}$. Let $\phi$ be a half twist that swaps $\diamond$ and $f^{q-1}(C_{1})$ (which could be $C_{1}=f^{0}(C_{1})$). Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$ (in this case not all points of $M$ are vertices of the portrait), and the mapping class $\phi$. Then the portrait of $f\phi$ has two components: $C_{1}$ is in a cyclic component of length q1q-1 and $C_{2}$ is in a component that is the same as the forward orbit of $C_{2}$ under $f$. In particular, $C_{1}$ is cyclic under $f\phi$ and $C_{2}$ is strictly pre-periodic under $f\phi$.
@@
[…]
    Figure 2. Function 1(4) may change all portrait features of $\Gamma_{f}$